OpenStudy (ineedhelpfast12):

Solve the linear system using elimination. 10x-9y=46 -2x+3y=10

OpenStudy (anonymous):

multiply the second equation all the way across by 3, then add the two together and the y terms will go bye bye

OpenStudy (ineedhelpfast12):

so x=4?

OpenStudy (anonymous):

yes

OpenStudy (ineedhelpfast12):

but 10(4)-9y wont equal 46

OpenStudy (anonymous):

you have to solve for y

OpenStudy (ineedhelpfast12):

ok so 40-9y=46. 9y=6?

OpenStudy (anonymous):

oh i was wrong it is not x = 4

OpenStudy (anonymous):

\[4x=76\]

OpenStudy (anonymous):

when you add

OpenStudy (anonymous):

so \(x=19\) not \(4\)

OpenStudy (anonymous):

sorry about that

OpenStudy (ineedhelpfast12):

so 10x-8y=46 -6x+9y=30

OpenStudy (anonymous):

\[10x-9y=46\\ -6x+9y=30\]

OpenStudy (ineedhelpfast12):

y=16?

OpenStudy (anonymous):

\[4x=76\\ x=19\]

OpenStudy (ineedhelpfast12):

therefore, if x=19, y=16

OpenStudy (anonymous):

yes

OpenStudy (ineedhelpfast12):

Thank you! Can you please help me with a couple more?

OpenStudy (anonymous):

sure if i can

OpenStudy (ineedhelpfast12):

8x-5y=11 4x-3y=5

OpenStudy (anonymous):

this one is ok too, multipliy the second one by \(-2\) then add

OpenStudy (ineedhelpfast12):

why -2?

OpenStudy (anonymous):

because then you will have \(8x\) and \(-8x\)

OpenStudy (anonymous):

this is actually what you are supposed to be learning, what to multiply by to eliminate one of the variables

OpenStudy (ineedhelpfast12):

so y=1, then I solve for x

OpenStudy (ineedhelpfast12):

x=2?

OpenStudy (anonymous):

yes

OpenStudy (ineedhelpfast12):

how about this one. 11x-20y=28 3x=4y=36

OpenStudy (anonymous):

too many equal signs

OpenStudy (ineedhelpfast12):

whoops! 3x+4y=36

OpenStudy (anonymous):

ok now it is your turn figure out what to multiply the second one by to get rid of one of the variables

OpenStudy (ineedhelpfast12):

5?

OpenStudy (anonymous):

yes

OpenStudy (ineedhelpfast12):

so 11x-20y=28 15x+20y=36

OpenStudy (anonymous):

right and now if you add the y terms go away that is what you are learning here

OpenStudy (ineedhelpfast12):

so 15x=180

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

i think there is a mistake somewhere did you multiply both sides by 5?

OpenStudy (ineedhelpfast12):

both equation 1 and 2 no, but I multiplied 3x, 4y, and 36 by 5

OpenStudy (anonymous):

so you have \[11x-20y=28\\ 15x+20y=180 \]

OpenStudy (ineedhelpfast12):

yes

OpenStudy (anonymous):

add and get what?

OpenStudy (ineedhelpfast12):

26x=208

OpenStudy (ineedhelpfast12):

so does x=8 and y=3?

OpenStudy (anonymous):

yes

OpenStudy (ineedhelpfast12):

this next one is a word problem

OpenStudy (anonymous):

great

OpenStudy (ineedhelpfast12):

A library is having a book sale to raise money. Hardcover books cost $4 each and paperback books cost $2 each. A person spends $26 dollars for 8 books. How many hardcover books did she purchase?

OpenStudy (anonymous):

you can do this a bunch of ways, but i guess they want you to write two linear equation the total is 8 books so if x is the hard cover, y the paper back then \[x+y=8\]

OpenStudy (anonymous):

the total cost is 26 so using the above variables you have \[4x+2y=26\] solve that system

OpenStudy (ineedhelpfast12):

x=5 and y=3?

OpenStudy (anonymous):

yeah an easy check

OpenStudy (anonymous):

\[5+3=8\\ 4\times 5+2\times 3=26\]

OpenStudy (ineedhelpfast12):

What about this one? A website allows users to download individual songs or an entire album. All individual songs cost the same to download, and all album cost the same to download./ Ryan pays $14.94 to download 5 individual songs and 1 album. Seth pays $22.95 to download 3 individual songs and 2 albums. How much does the website charge to download a songs? An entire album?

OpenStudy (anonymous):

you want to try the two equations

OpenStudy (ineedhelpfast12):

how do I get them though?

OpenStudy (anonymous):

i will do the first, you can do the second

OpenStudy (anonymous):

if x is songs, y is lps then \[5x+y=14.95\]

OpenStudy (ineedhelpfast12):

so 3x+2y=$22.95

OpenStudy (anonymous):

yup

OpenStudy (ineedhelpfast12):

how do I solve it

OpenStudy (anonymous):

\[3x+2y=22.95\\ 5x+y=14.95\]go ahead and figure out what to multiply the second one by to eliminate the y

OpenStudy (ineedhelpfast12):

2

OpenStudy (anonymous):

actually \(-2\)